.Consider a schedule with 3 transactions with 2, 3, and 4 operations respectively. Calculate the possible

number of

c) Serial schedules

d) Non serial schedules


1
Expert's answer
2021-11-09T17:46:29-0500
= (1 + 2 + 3)! / (1! + 2! + 3!) = 6! / 9 = 120

Number of Serial Schedules,
= 3! = 6
Number of Non-Serial Schedules,
= Total Number of Schedules - Number of Serial Schedules 
= 120 - 6 
= 114